Circular No. 6838

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SAX J1810.8–2609 P. Ubertini, Istituto di Astrofisica Spaziale (IAS), CNR, Rome; J. in ’t

Zand, Space Research Organization of the Netherlands, Utrecht; A. Tesseri and D. Ricci, BeppoSAX Science Operation Center, Rome; and L. Piro, IAS, report on the detection of a new weak x-ray transient in the galactic bulge with the Wide Field Camera 2 onboard BeppoSAX. The source has been detected since Mar. 10.18 UT at a steady level of 15 mCrab (2–9 keV), at α = 18^{h}10^{m}46^{s}, δ = −26^{o}09′.1 (equinox 2000.0; preliminary error radius 3′). An x-ray burst of duration 45 s was detected from this source on Mar. 11.06634. The peak intensity is ∼ 1.6 Crab. The only catalogued source in the error box is IRAS 18077–2609, located 1′.7 away. A BeppoSAX target-of-opportunity observation is in progress. Observations at other wavelengths are encouraged.

SUPERNOVA 1998S IN NGC 3877 Visual magnitude estimates: Mar. 11.79 UT, 12.0, (K. Hornoch, Lele-

kovice, Czech Republic; comparison stars from GSC); 11.83, 12.0 (L. Brat, Brno, Czech Republic); 11.83, 12.3 (L. Novak, Brno, Czech Republic).

COMET C/1997 J2 (MEUNIER-DUPOUY) Ephemeris extension to IAUC 6801 (orbital elements MPC 31069):

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 28 21^{h}48.^{m}02 +31^{o}43.^{′}9 3.647 3.056 47.^{o}1 13.^{o}8 11.2
Apr. 7 21 58.77 +31 26.8 3.617 3.063 49.5 14.4 11.2
17 22 08.17 +31 13.3 3.570 3.073 53.0 15.1 11.1
27 22 16.13 +31 01.5 3.507 3.087 57.6 16.0 11.1
May 7 22 22.54 +30 48.9 3.427 3.103 63.1 16.8 11.1
17 22 27.26 +30 32.8 3.333 3.122 69.3 17.6 11.1
27 22 30.14 +30 10.2 3.227 3.145 76.3 18.2 11.0
June 6 22 31.00 +29 37.1 3.111 3.170 84.0 18.6 11.0
16 22 29.70 +28 49.2 2.990 3.198 92.4 18.5 10.9

Total magnitude estimates, visual unless otherwise noted: Feb. 1.72 UT, 10.6 (M. Reszelski, Szamotuly, Poland, 0.25-m reflector); 5.18, 11.1 (K. Hornoch, Lelekovice, Czech Rep., 0.35-m refl.); Mar. 10.17, 12.3 (H. Mikuž, Crni Vrh Observatory, Slovenia, 0.36-m reflector + CCD + V filter).

1998 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green